OpenStudy (anonymous):

find the laplace transform of tcos(2t)

hartnn (hartnn):

do u know the L.T of cos 2t ?

OpenStudy (anonymous):

yes so its s/(s^2 + 4)

OpenStudy (anonymous):

so do i have to take the integral of the function first? or...

hartnn (hartnn):

now u need to use \(\large L[t^nf(t)]=(-1)^n\frac{d^n}{ds^n}[F(s)]\) did u know this ?

OpenStudy (anonymous):

no i dont take the integral theres got to be an easier way

OpenStudy (anonymous):

okay but then what is F(s) in this case?

hartnn (hartnn):

F(s) is the Laplace transfom of f(t) in this case, f(t) is cos 2t and F(s) is its L.T

OpenStudy (anonymous):

so it would be -1 * d/ds * F(s)

hartnn (hartnn):

thats why i first asked whether u know L.T of cos 2t

hartnn (hartnn):

YUP

OpenStudy (anonymous):

wait so why isnt f(t) tcos(2t) and is cos(2t) haha

hartnn (hartnn):

because t^1 is multiplied by cos 2t so, according to t^n f(t) f(t) is cos 2t and n=1

OpenStudy (anonymous):

oh my gosh thats why im messing up thank you so much hartnn :)

hartnn (hartnn):

welcome ^_^ i hope you can finish now....

OpenStudy (anonymous):

got it!:) thanks again